OpenStudy (anonymous):

solve sinx+2sinxcosx=0 for 0 degrees ≤ x ≤ 180 degrees

OpenStudy (aravindg):

sin x+2sinx cos x=sin x(1+2cos x )=0 implies sinx=0 or 1+2cos x=0 solve :)

OpenStudy (anonymous):

Thank you c:

OpenStudy (anonymous):

so sin0(1+2cos0) = 0?

OpenStudy (aravindg):

not that way solve each separately sin x=0 x=? 1+2cos x =0 x= ?

OpenStudy (aravindg):

note that you need to report a general value in 0 degrees ≤ x ≤ 180 degrees

OpenStudy (anonymous):

I don't get it :l

OpenStudy (aravindg):

ok i will do first one sin x=0 means x=0 or pi

OpenStudy (anonymous):

oh >_<

OpenStudy (aravindg):

can u do 2nd one ?

OpenStudy (anonymous):

1+2cos(pi)?

OpenStudy (aravindg):

first solve for cos x

OpenStudy (anonymous):

Would it not be pi? :l

OpenStudy (aravindg):

1+2cos x=0 cos x= ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

so 1+2cosx=3?

OpenStudy (aravindg):

............

OpenStudy (anonymous):

No? >_<

OpenStudy (anonymous):

I'm not good at these >_<

OpenStudy (aravindg):

1+2cos x=0 cos x=?

OpenStudy (anonymous):

Ohhhh 3 cuhs all of that equals 0?

OpenStudy (anonymous):

cos x = 3 I mean

OpenStudy (aravindg):

???

OpenStudy (anonymous):

>_<

OpenStudy (aravindg):

assume cos x=y 1+2y=0 y= ?

OpenStudy (anonymous):

-1/2?

OpenStudy (aravindg):

exactly !! well done :)

OpenStudy (aravindg):

finally :P

OpenStudy (anonymous):

lmao sorry I didn't get it cx

OpenStudy (anonymous):

Well at first cx

OpenStudy (anonymous):

So how would my final answer look? :l

OpenStudy (aravindg):

COS X=-1/2 x= 120 degree or pi-pi/3

OpenStudy (anonymous):

oh okai c: thank you c:

OpenStudy (aravindg):

yw :)